OpenStudy (anonymous):

how to integrate pie*(cosx^2)^2 dx ?thx :)

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=integrate+%5C%5BPi%5D*%28cosx%5E2%29%5E2 Click on "Show steps"

OpenStudy (jamesj):

Do you want the integral of \( \pi \cos^2(x^2) \) or the integral of \( \pi (\cos^2 x)^2 \)?

OpenStudy (jamesj):

The answer to the second integral is what robtobey has given you. The answer to the first question is that there is no expression for the integral of the function in terms of elementary functions.